OpenStudy (anonymous):

If f'(x)=-f(x) and f(1)=1, what's f(x)? (Answer: e^(1-x))

OpenStudy (owlcoffee):

That is correct.

OpenStudy (anonymous):

But I don't know how to do the problem.

OpenStudy (kc_kennylau):

y'=-y, y'+y=0 Auxiliary equation: r+1=0, r=-1 General solution: y=ae^(-x) f(1)=1: ae^(-1)=1, a=e. Therefore solution: y=e*e^(-x)=e^(1-x)

OpenStudy (anonymous):

Thank you.

OpenStudy (kc_kennylau):

no problem :)